OpenStudy (anonymous):

3t+9t=60 what is this answer? Please!!!

OpenStudy (anonymous):

5

OpenStudy (anonymous):

5... 3t+9t=12t... so 12t=60... 60/12=5

OpenStudy (anonymous):

Thank you all so much! Having trouble.

OpenStudy (luigi0210):

Welcome to Openstudy. And no giving away answers guys.

OpenStudy (anonymous):

3w +4W-2=12 Can someone help me solve this problem?